OpenStudy (anonymous):

find the exact value of the expression sin(5pi/12)

OpenStudy (anonymous):

.02284431908 i used a calculator

OpenStudy (mertsj):

\[\sin \frac{5\pi}{12}=\sin (\frac{3\pi}{12}+\frac{2\pi }{12})=\sin (\frac{\pi}{4}+\frac{\pi}{6})\]

OpenStudy (mertsj):

Now use this formula: \[\sin (a+b)=\sin a \cos b+\cos a \sin b\]